OpenStudy (anonymous):

f(x)=e^(1/x)...i need help explaining the graph and how the exponent affects the graph's different portions so when x<0, 00.

OpenStudy (anonymous):

type the equation in on wolframalpha.com and u can see the graph. basically when x<0 the function value is getting closer and closer to one as it reaches negative infinity, but it never =1. between 0 and 1, the function gets closer and closer to positive infinity as x gets closer to 0. And when x is greater then zero the function gets closer and closer to 1 as the x value goes to infinity.

OpenStudy (anonymous):

that what your looking for bro?

OpenStudy (anonymous):

ok so let me see if I understand, so when x<0 the graph is closer to which part of the function 1/x? :/ i think thats how she sort of wants to explain it like which part of the function it is closest to

OpenStudy (anonymous):

well hm are there even two parts to the function its not like its e^(x) + 1/x where the two parts would be e^(x) then 1/x..

OpenStudy (anonymous):

when x is less then zero the funtion (f(x)) gets closer to negative infinity so basically in terms of the exponent, as the exponent decreases (as in when x gets smaller and smaller as x<0 indicates) then function f(x), approaches one. So like if x is-50, f(x) will be like .87 for example,and when x is -2000 f(x) woll be like .99, and -50000 f(x) will be like .9999999

OpenStudy (anonymous):

theres 1 part to the function

OpenStudy (anonymous):

or nvm what were you asking?

OpenStudy (anonymous):

so the exponent dominates basically so to speak when x approaches negative infinity because..?

OpenStudy (anonymous):

i dont understand your question.k, lets back track. did u read what i wrote about f(x) going to like .9999?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok did u not understnad it?

OpenStudy (anonymous):

think of 1/x. if x is small (like -500000), what will (1/x) =???.. and.......... if x isbig (like 50000000), what will (1/x) =??

OpenStudy (anonymous):

hm not sure

OpenStudy (anonymous):

well when you have 1/5000 you are putting a larger number on the denomenator so you get a very small fraction? and then if you have 1/-500 you get..

OpenStudy (anonymous):

you get a really small negative fraction?

OpenStudy (anonymous):

small fraction again just negative

OpenStudy (anonymous):

yes. so when the x value gets big. it approaces 0 right.

OpenStudy (anonymous):

and when it gets small it a pproaches negative infinity

OpenStudy (anonymous):

hm okay just trying to figure out how to say all of this hmm lol

OpenStudy (anonymous):

so that answers part of your question

OpenStudy (anonymous):

x<0 ... thats where we have negative numbers like -2, -50, -500000. o as we keep mmaking x smaller in x<0 , the overall number for (1/x) get super small too...and thus approaches negative infinity. u gotta be able to grasp that .

OpenStudy (anonymous):

no i do now okay, and how come in between 0 and 1 it approaches 0

OpenStudy (anonymous):

well when x=1 eaxctly f(x) just equals e^1 right?

OpenStudy (anonymous):

e^(1/x) ....e^(1/1) = e^(1) ....1/1 =1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and that is just some number like 2 .7 right. so thats one edge of the boundary. the other edge is 0. cuz it wants x between 0 and 1.

OpenStudy (anonymous):

no wait e^(1) is just e not 1 e^(1) is not 1 its e

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ya e^1=e

OpenStudy (anonymous):

when x=0 it says f(x) is undefined..shouldnt it be 1 because any exponent of 0 =1

OpenStudy (anonymous):

no. think about it1/x....if x=0, u have 1/0- thats what we call that undefined and you'll get an error on your calculator cuz u cant devide by 0.

OpenStudy (anonymous):

5/0. no such thing. 7/0. no such thing. elephant/0. no such thing.

OpenStudy (anonymous):

its a hard one to grasp just accept it. 0 is nothing so if you divide something by nothing u get nothing.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ight moving on. so between 0 and 1 we have something that equals a number between (undefined and 2.7 right??)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

if we back it upa bit. and instead of doing1/0. we do 1/say 0.000001 (which is really close to 0. we get a HUGE number for (1/x). right? u get that.

OpenStudy (anonymous):

therefore we can say thaas x gets closer to 0 , from 1 (between 0 and 1) the 1/x statement reaches INFINITY. got it?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

cuz HUGE number is close to Infinity....

OpenStudy (anonymous):

can we do f(x)=x^(1/x) (same thing but x instead of e before the exponent)

OpenStudy (anonymous):

no my man. e is a constant. its not x.

OpenStudy (anonymous):

no no the next one i have to do..is f(x)=x^(1/x)

OpenStudy (anonymous):

oh. ythats a little different put that into wolframalpha and u can see the graph you'd get. ight i gotta peace, i got some lawn work to take care of. peace man hope that was helpful!!

OpenStudy (anonymous):

thankssssssssss